Builder Products, Inc., uses the weighted-average method in its process costing system. It manufactures a caulking compound that goes through three processing stages prior to completion. Information on work in the first department, Cooking, is given below for May:

Production data:
Pounds in process, May 1; materials 100% complete;
conversion 90% complete
83,000
Pounds started into production during May 480,000
Pounds completed and transferred out ?
Pounds in process, May 31; materials 75% complete;
conversion 25% complete
43,000
Cost data:
Work in process inventory, May 1:
Materials cost $ 128,300
Conversion cost $ 53,900
Cost added during May:
Materials cost $ 666,940
Conversion cost $ 296,395

Required:

1. Compute the equivalent units of production for materials and conversion for May.

2. Compute the cost per equivalent unit for materials and conversion for May.

3. Compute the cost of ending work in process inventory for materials, conversion, and in total for May.

4. Compute the cost of units transferred out to the next department for materials, conversion, and in total for May.

5. Prepare a cost reconciliation report for May.
Answer #1

1) Equivalent unit

Material Conversion
Unit transferred out 83000+480000-43000 = 520000 520000
Ending WIP 43000*75% = 32250 10750
Total 552250 530750

2) Cost per equivalent unit

Material Conversion
Beginning WIP 128300 53900
Cost added 666940 296395
Total Cost 795240 350295
Equivalent unit 552250 530750
Cost per equivalent unit 1.44 0.66

3) Cost of ending WIP

Material Conversion Total
Cost of ending WIP 32250*1.44 = 46440 10750*.66 = 7095 53535

4) Cost of Unit transferred out

Material Conversion Total
Cost of Unit transferred out 520000*1.44 = 748800 520000*.66 = 343200 1092000

5) Cost reconciliation

Cost to be accounted for
Beginning WIP 182200
Cost added 963335
Total cost to be accounted for 1145535
Cost accounted as
Cost of unit transferred out 1092000
Ending WIP 53535
Total cost accounted for 1145535